OpenStudy (chris215):

verify the identity. cot(x-pi/2) = -tan x

OpenStudy (misty1212):

HI!!

OpenStudy (chris215):

hi!

OpenStudy (misty1212):

what does "verify" mean in this instance? my guess is it means use the subtraction angle formula for cotangent but i could be wrong

OpenStudy (chris215):

im not sure

OpenStudy (misty1212):

i don't know it, so i just looked it up it is \[\cot(A-B)=\frac{\cot(A)\cot(B)+1}{\cot(B)-\cot(A)}\] sort of annoying

OpenStudy (misty1212):

actually now it is easy, since \[\cot(\frac{\pi}{2})=0\]

OpenStudy (misty1212):

\[\cot(A-B)=\frac{\cot(A)\cot(B)+1}{\cot(B)-\cot(A)}\] \[\cot(x-\frac{\pi}{2})=\frac{\cot(x)\cot(\frac{\pi}{2})+1}{\cot(\frac{\pi}{2})-\cot(x)}\]

OpenStudy (misty1212):

you get \[-\frac{1}{\cot(x)}\] right away

OpenStudy (chris215):

oh ok thanks so much :)

OpenStudy (chris215):

actually I think I'm supposed to show how cot (x-pi/2)= -tan x

OpenStudy (matt101):

Keep in mind that cot and tan are inverse functions, so: -1/cotx = -tanx